OpenStudy (anonymous):

Find the equilibrium point of the demand and supply equations. (The equilibrium point is the price p and number of units x that satisfy both the demand and supply equations.) Demand Supply p = 133 – 0.09x p = 0.9x – 560 A. (700, 70) B. (400, 97) C. (857, 211.3) D. (600, 79) E. inconsistent

OpenStudy (radar):

A

OpenStudy (radar):

If you want to know why let me know/

OpenStudy (anonymous):

Yes I would please! This is a practice problem and I probably will have more on my test ...

OpenStudy (radar):

P=P so then these two equations are equal for equlibrum: 133-0.09x=0.9x-560 Solve for x by consolidating/simplifying getting .99x=693 or x=700 substitute the 700 in your P=0.9(700) - 560=630-560=70, P=70

OpenStudy (anonymous):

Awesome thank you!

OpenStudy (radar):

No problemo